# via Dodecaneso 35, 16146, Genova, Italy

### v1

@@ 6.4 A linearization attack for $s=0$
First, we consider a Pesto primitive with $t=1$ (so $x=(x_{1})$), $s=0$ and such that $U(x,y)=U(x_{1},y)$ does not have the quadratic term $x_{1}^{2}$. In this case we have $U(x_{1},y)=x_{1}\alpha(y)+\beta(y)$, with $\alpha,\beta:{\mathbb{F}}_{q}^{n-1}\rightarrow{\mathbb{F}}_{q}^{m-1}$ affine maps. Hence, for $G(x_{1},y)=(c_{T},c_{U})$ it holds that
@@
This same analysis can be generalized to the case $t>1$ and $s=0$. Indeed, the map $U$ results to have terms of the form $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{i}$, $y_{j}$. So, given $G(x,y)=(c_{T},c_{U})$ we have $m-t$ quadratic equations in $y,c_{T},c_{U}$, with the variables in $y$ appear only in degree 1. Hence, also from $G_{pub}$ it is possible to recover an equation $B(\mathfrak{i},\mathfrak{o})=0$ which is quadratic but with the variables of the input appearing only in degree 1. Hence the same attack can be performed. For this reason, we choose $s>0$ in Definition 4.2.
@@ 6.5 Known attacks on Oil and Vinegar systems
Since the second part of the system $G$ is formed by polynomials of degree up to 4, we believe that the usual OV attacks cannot easily be applied to the whole map $G_{pub}$.
@@ 6.6 Algebraic attack with Gröbner bases
In this section, we consider algebraic attacks using Gröbner bases. The scenario is that an attacker wants to forge a signature, hence to find a preimage of a random element. We consider a random value $w\in{\mathbb{F}}_{q}^{m}$ (the hash of a document), the goal is to find $v\in{\mathbb{F}}_{q}^{n}$ such that $G_{pub}(v)=w$. We can do this by finding a Gröbner basis of the polynomial system $G_{pub}$ by the usual strategy (see e.g. [CG21b]) with linear-algebra-based-algorithms such as F4 [Fau99], F5 [Fau02], XL [CKPS00], etc. The complexity of these algorithms is bounded from above by
$O\left(\binom{n+{\mathrm{sd}}(G_{pub})}{n}^{\omega}\right)$
where $n$ is the number of variables, ${\mathrm{sd}}(G_{pub})$ is the solving degree, and $32<\omega<$ 3. In a nutshell, the solving degree represents the highest degree of polynomials that need to be considered during the process of solving the system [CG22]. In order to estimate this for the system $G_{pub}$ of Pesto, we performed some computational experiments with MAGMA software [BCP97] for different values of the parameters $n,m,t,s,q$. We also computed the Castelnuovo–Mumford regularity of the homogenized system $G_{pub}^{h}$, which gives an upper bound for the solving degree [CG21b]. The results are summarized in the following table.
$n$ $m$ $t$ $s$ ${\mathrm{sd}}(G_{pub})$ $\mathrm{reg}(G_{pub}^{h})$
$q=5$ 5 4 2 1 4.54 7.6
$q=2^{6}$ 5 4 2 1 4.06 7.96
$q=3761$ 5 4 2 1 4 8
$q=5$ 6 5 2 2 7.64 8.74
$q=2^{6}$ 6 5 2 2 7.94 9
$q=3761$ 6 5 2 2 8 9
Table 3: Mean values of solving degree and regularity for 50 randomly generated systems
For higher parameters, we were not able to compute the above values. Therefore, we added to $G_{pub}$ the field equations $x_{i}^{q}-x_{i}=0$, $y_{i}^{q}-y_{i}=0$ and then we computed the solving degree for this larger set of equations. Clearly, this strategy is not efficient for large values of $q$, hence we only considered $q=5$. As for the previous experiments, we considered some different randomly generated systems $G_{pub}$ and then we display the mean value for ${\mathrm{sd}}(G_{pub})$, see Table 4.
@@
Table 4: Mean value of the solving degree for $25$ randomly generated systems $G_{pub}$ over $\mathbb{F}_{5}$
7 Conclusions
@@
In this work, we propose applying a CCZ transformation in the construction of a multivariate scheme, instead of the usual affine transformation. This has the advantage of hiding linear relations between the input and output that would occur with an affine transformation alone. However, this approach may increase the public key size.
@@
With this work, we hope to build a fruitful bridge between the areas of cryptographic Boolean functions and Multivariate Cryptography. Since both areas work with functions/polynomials defined over finite fields, we believe that many techniques used in one area could be studied and applied in the other. We see significant potential for further exploration in this direction.

### v2

@@ 6.4 A linearization attack for $s=0$
First, we consider a Pesto scheme with $t=1$ (so $x=(x_{1})$), $s=0$ and such that $U(x,y)=U(x_{1},y)$ does not have the quadratic term $x_{1}^{2}$. In this case we have $U(x_{1},y)=x_{1}\alpha(y)+\beta(y)$, with $\alpha,\beta:{\mathbb{F}}_{q}^{n-1}\rightarrow{\mathbb{F}}_{q}^{m-1}$ affine maps. Hence, for $G(x_{1},y)=(c_{T},c_{U})$ it holds that
@@
This same analysis can be generalized to the case $t>1$ and $s=0$. Indeed, the map $U$ results to have terms of the form $x_{i}x_{j}$, $x_{i}y_{j}$, $x_{i}$, $y_{j}$. So, given $G(x,y)=(c_{T},c_{U})$ we have $m-t$ quadratic equations in $y,c_{T},c_{U}$, with the variables in $y$ appearing only in degree 1. Hence, also from $G_{pub}$ it is possible to recover an equation $B(\mathfrak{i},\mathfrak{o})=0$ which is quadratic but with the variables of the input appearing only in degree 1. Hence, the same attack can be performed. On the other hand, if $s>0$, the attack in its current form cannot be performed because the resulting relation $c_{U}=U(c_{T},y)$ will not be linear in the input variables $y$. For this reason, we choose $s>0$ in Definition 4.2.
@@ 6.5 Known attacks on Oil and Vinegar systems
On the other hand, since the second part of the system $G$ is formed by polynomials of degree up to 4, we believe that the usual OV attacks cannot be applied directly to the whole map $G_{pub}$.
@@ 6.6 Algebraic attack with Gröbner bases
In this section, we consider algebraic attacks using Gröbner bases. The scenario is that an attacker wants to forge a signature, hence to find a preimage of a random element. We consider a random value $w\in{\mathbb{F}}_{q}^{m}$ (the hash of a document), the goal is to find $v\in{\mathbb{F}}_{q}^{n}$ such that $G_{pub}(v)=w$. We can do this by finding a Gröbner basis of the polynomial system $G_{pub}$ by the usual strategy (see e.g. [CG21b]) with linear-algebra-based-algorithms such as F4 [Fau99], F5 [Fau02], XL [CKPS00], etc. For a large class of these algorithms, the complexity can be bounded from above by
$O\left(\binom{n+{\mathrm{sd}}(G_{pub})}{n}^{\omega}\right)$ (5)
where $n$ is the number of variables, ${\mathrm{sd}}(G_{pub})$ is the solving degree, and $32<\omega<$ 3. In a nutshell, the solving degree represents the highest degree of polynomials that need to be considered during the process of solving the system [CG22]. In order to estimate this for the system $G_{pub}$ of Pesto, we performed some computational experiments with MAGMA software [BCP97] for different values of the parameters $n,m,t,s,q$. For each set of parameters, we randomly generated $50$ systems and report the mean value of the solving degree. The results are summarized in Table 2.
@@
Table 2: Mean values of the solving degree for 50 randomly generated systems.
Remark 6.7.
We also conducted similar experiments for smaller field sizes, specifically for $q\leq 5$. As expected, in this scenario, it is advantageous to add the field equations $x_{i}^{q}-x_{i}=0$ and $y_{i}^{q}-y_{i}=0$ to the system $G_{pub}$ when computing the Gröbner basis. In several cases, this approach results in a lower solving degree compared to those reported in Table 2. However, this strategy becomes inefficient for larger values of $q$. Therefore, we recommend implementing the scheme over larger fields, specifically choosing $q\geq 2^{6}$. Moreover, since our goal is to analyse how the solving degree of our scheme behaves for different values of $q$ (including large $q$), we believe that studying the system without the field equations will offer better insight into how the solving degree scales.
6.7 Parameter proposals
Taking into account all potential attacks explored so far, as well as the key size analysis in §5, we propose specific parameter choices for the signature scheme Pesto from Definition 4.2. Based on our analysis, we set $t\approx\frac{n}{3}$ and $s=n-m=2$. Under these assumptions, the previously outlined structural attacks do not appear to be applicable. In this scenario, the most effective attack remains the direct algebraic attack using Gröbner bases. To mitigate this, we choose a large field size ($q=2^{6}$) to make adding field equations to the system computationally infeasible. As shown in Table 2, the solving degree of the public system appears to grow rapidly and linearly with $n$, likely due to the presence of degree 4 equations in $G_{pub}$. Based on this observation, we conjecture possible estimates for the solving degree with larger parameters and use them to assess the complexity of a Gröbner basis attack, applying (5) with $\omega=2.3$.
We then select the remaining parameters $(n,m,t)$ to align with NIST security levels for post-quantum signature schemes and compute the corresponding public and secret key sizes using the formulas from Remark 6.1. For comparison, we evaluate these key sizes against those of the UOV signature scheme [BCD+] (without compression), which has been submitted to NIST’s Call for Additional Digital Signature Schemes in the PQC Standardization Process. The results are summarized in Table 3.
Security level Parameters $(\mathbb{F}_{q},n,m,t)$ Estimated ${\mathrm{sd}}(G_{pub})$ Pesto PK size (MB) Pesto SK size (MB) UOV public key size (MB) UOV secret key size (MB)
NIST I $(\mathbb{F}_{2^{6}},27,25,10)$ $\geq 33$ 357 5.5 278.5 237.9
NIST III $(\mathbb{F}_{2^{6}},40,38,14)$ $\geq 48$ 2453 16.5 1225.4 1044.3
NIST V $(\mathbb{F}_{2^{6}},57,55,20)$ $\geq 59$ 13725 43.8 2869.4 2436.7
Table 3: Parameter proposals for the signature scheme Pesto of Definition 4.2.
Remark 6.8.
We would like to emphasize that the parameters in Table 3 are merely indicative of how the key sizes of Pesto might grow. In particular, they highlight that there appears to be a clear advantage in the secret key size, while the public key size increases rapidly, which could make the scheme more suitable for lower security levels. However, at this stage of the proposal, it remains uncertain whether our assumptions about how the solving degree scales with the parameters are correct. We therefore encourage further cryptanalysis and additional study of the scheme to better assess its security.
7 Challenges and open problems
@@
In this work, we propose applying a CCZ transformation in the construction of a multivariate scheme, as an alternative to the usual affine transformation. This offers the advantage of concealing linear relationships between the input and output that would typically arise with an affine transformation alone. To better evaluate the stability and feasibility of this approach, we have identified a series of open problems and questions relevant to future cryptanalysis.
  1. 1.
    How does the solving degree grow with respect to the input parameters? Due to the limited set of experiments we were able to conduct, predicting the behavior of the solving degree of the public polynomial system remains challenging. Based on our experiments, it appears that the solving degree scales linearly with the number of variables $n$. However, further investigation and, hopefully, a mathematical proof are desirable.
  2. 2.
    Is it possible to combine the CCZ construction with other known multivariate primitives? More specifically, can the map $U$ be chosen using HFE polynomials? How does the resulting scheme compare to the Pesto scheme of Definition 4.2?
  3. 3.
    None of the attacks presented in Section 6 appear to fully compromise the scheme. However, we believe that further cryptanalysis is necessary to better assess the security of the scheme. In particular, can classical attacks on Oil and Vinegar schemes be adapted to the Pesto scheme?
  4. 4.
    The public key consists of degree 4 polynomials and, as a result, appears to be quite large. At present, this is the main practical disadvantage of the scheme. Could techniques used for key size reduction in OV schemes, such as the QR technique from [FIKT21], be adapted to reduce the key size?
@@
As the previous problems highlight, we see significant potential for further exploration in this direction, which we hope could build a fruitful bridge between the areas of cryptographic Boolean functions and multivariate cryptography. Since both fields involve functions and polynomials defined over finite fields, we believe that many techniques used in one area could be studied and applied in the other.